# Knockoffs for low dimensions: changing the nominal level post-hoc to gain power while controlling the FDR

### v1

@@ 5.2 Evaluation measures
For each model and each signal amplitude setting, the dataset $\mathcal{D}=(\mathbf{X},Y)$ is generated independently 100 times. For each draw of the dataset, we apply the knockoffs procedures that we are comparing. For each model, each signal amplitude, and each of the two methods, we measure performance in two different ways. First, we estimate the power and FDR as
$\textnormal{Power}=\frac{1}{100}\sum_{d=1}^{100}\frac{|R_{d}\cap I_{1}|}{|I_{1}|}\quad\textnormal{and}\quad\textnormal{FDR}=\frac{1}{100}\sum_{d=1}^{100}\frac{|R_{d}\cap I_{0}|}{|R_{d}|\vee 1},$ (32)
where $R_{d}$ denotes the rejected set for the $d$-th dataset $\mathcal{D}_{d}$.
@@
$\textnormal{Average ratio FDP/}\alpha=\frac{1}{100}\sum_{d=1}^{100}\frac{|R_{d}\cap I_{0}|}{|R_{d}|\vee 1}\frac{1}{\alpha_{d}},$ (33)
@@ 5.3 Results
5.3.1 Comparing with the original knockoffs in low-dimensional settings
@@
In this section we illustrate the performance of our proposed post-hoc knockoff (presented in Algorithm 1) and compare to the original knockoff procedure from Candès et al. [4]. The details of the simulation setup are presented in Section 5.1. Through the simulation we set the $\alpha^{\textnormal{kn}}$ level to $0.20$ FDR, i.e., this is the nominal level of the traditional knockoffs, and the initial significance level of our post-hoc knockoffs.
@@
In Figure 2, we evaluate the performance of the different methods as a function of $p_{\text{relevant}}$, with signal strength held fixed. When the number of truly relevant variables is small, traditional methods suffer from very low power. In contrast, post-hoc knockoffs offer the flexibility to substantially increase power, with paying some price in the probability of committing a Type-I error. For $p_{\text{relevant}}>3$, the increase in Type-I error is small relative to the power gain. Importantly, across nearly all settings, the empirical FDR remains close to the nominal target level $\alpha$, comparable to that of the original knockoff procedure. Furthermore, when power is comparable across methods (e.g., Gaussian setting for $p_{\text{relevant}}>6$), the post-hoc approach can achieve stricter error control, resulting in effective $\alpha$ values less than $0.20$.
@@
Similar behavior is observed when $p_{\text{relevant}}$ is held constant and the signal amplitude is varied, as shown in Figure 3. Post-hoc knockoffs consistently achieve higher power than traditional methods, especially at moderate signal strengths, with the price of increasing $\alpha$ above $0.20$. For high signal amplitudes, our method achieves high power, and at the same time, it can provide more stringent control, i.e. $\alpha<0.20.$ Interestingly, the empirical FDR remains below the nominal $\alpha$ level in all scenarios.
Gaussian
Logistic
[…]
Gaussian
Logistic
[…]
5.3.2 Comparing with the derandomized knockoffs in high-dimensional settings
@@
In this section, we illustrate the performance of our proposed derandomized post-hoc knockoff procedure (presented in Algorithm 2) and compare it with the original derandomized knockoff procedure from Ren and Barber [21]. The details of the simulation setup are presented in Section 5.1. For the nominal levels, we adopt the configurations from Ren and Barber [21]: we set the FDR level $\alpha=0.20$, which corresponds to the $\alpha_{\textnormal{ebh}}$ level of the derandomized knockoffs and serves as the initial significance level in our post-hoc derandomized knockoffs procedure, while in both methods we set $\alpha^{\textnormal{kn}}=\frac{\alpha_{\textnormal{ebh}}}{2}.$
@@
Figure 4 presents the results under the Gaussian linear model, where the dimensionality is $p=800$. The last column, corresponding to $p_{\text{relevant}}=80$, reflects the setting considered in Ren and Barber [21]. In this regime, both methods exhibit comparable performance overall. However, our post-hoc procedure yields a modest power gain at lower signal amplitudes, albeit with a slight increase in the Type I error rate. This trade-off becomes more pronounced in sparser settings, where $p_{\text{relevant}}=20$ and $40$. Notably, in highly sparse scenarios, our method can also achieve a lower Type I error rate while maintaining power comparable to the original derandomized approach. For example, when $p_{\text{relevant}}=20$ (first column) and the signal amplitude exceeds $5$, our procedure attains similar power with improved Type I error control.
Figure 5 presents the results under the logistic regression model, where the dimensionality is $p=600$. The final column, corresponding to $p_{\text{relevant}}=60$, serves as the primary reference setting. In this regime, both methods demonstrate overall comparable performance. Nevertheless, our post-hoc procedure provides a moderate power improvement at lower signal amplitudes, accompanied by a slight increase in the Type I error rate. This trade-off becomes more pronounced in sparser configurations. Notably, in the most sparse regimes, the post-hoc method can also achieve a lower Type I error rate while maintaining power comparable to the original derandomized approach. In particular, for larger signal amplitudes, our method matches the power of the original procedure while offering improved Type I error control.
@@
Figure 4: Performance comparison between the proposed derandomized post-hoc knockoff procedure and the original derandomized knockoffs method of Ren and Barber [21] under a Gaussian linear model. Results are reported across signal amplitudes for varying sparsity levels, given by the number of relevant variables $p_{\text{relevant}}$, including the original setting $p_{\text{relevant}}=80$.
@@
Figure 5: Performance comparison between the proposed derandomized post-hoc knockoff procedure and the original derandomized knockoffs method of Ren and Barber [21] under a Logistic linear model. Results are reported across signal amplitudes for varying sparsity levels, given by the number of relevant variables $p_{\text{relevant}}$, including the original setting $p_{\text{relevant}}=60$.
6 A case study using synthetic clinical trial data
@@
To demonstrate the practical utility of our approach and the advantages of post-hoc knockoffs for analysing clinical trial data, we consider a synthetic dataset designed to mimic a real randomized controlled trial. The data were generated using the benchtm R package [30], using Scenario 1 with a continuous outcome and a sample size of $n=1000$. Each patient is described by $p=30$ baseline covariates, a dimensionality that reflects realistic clinical settings where tens of demographic, clinical, laboratory, and lifestyle variables are routinely collected and jointly analysed. The outcome for each patient was generated as a function of covariates $\mathbf{X}=(X_{1},X_{2},\dots,X_{30})$ and treatment assignment $T\in\{0,1\}$ according to
@@
where $\mathbb{I}(\cdot)$ denotes the indicator function, $s=2.32$ controls the baseline covariate effect, and $\beta_{0}=0.053$ and $\beta_{1}=0.383$ determine the average treatment effect and the additional effect in the specified subpopulation, respectively. These parameters were tuned to mimic real clinical trial data. A full description of the data-generating mechanism and simulation design is provided in Sun et al. [30].
In this setting, the target variables to be selected by the knockoff procedures are $X_{1}$, $X_{11}$ and the treatment indicator $T$.
We simulate 10 independent RCT datasets following the procedure described above, denoted by R1, …, R10. We examine which variables are selected by the two methods under a nominal FDR level of 0.20. Figure 6 displays the corresponding heatmaps of selected variables.
As shown in the left panel, the original knockoffs method selects five variables only in run R2, and selects none in the remaining nine runs. This implies that, 90% of the time, the method makes no discoveries, leading to frequent failures in detecting true signals. In contrast, the post-hoc method (right panel) selects variables $X_{1}$ and $X_{11}$, both true predictors, in all ten runs, and additionally selects the third relevant variable $T$ in 7 out of 10 runs. The average $\alpha^{ph}$ across the ten runs is 0.32. Moreover, the heatmap shows that the realized FDR is substantially below the nominal target, a trend that is consistently observed across the broader simulations in Section 5.

### v2

@@ 5.2 Evaluation measures
For each model and each signal amplitude setting, the dataset $\mathcal{D}=(\mathbf{X},Y)$ is generated independently D times, with D = 2000 for the low-dimensinal regime simulations (Sections 5.3.1, 5.3.4 and 5.3.2), and D = 200 for the high-dimensional (Section 5.3.3). For each draw of the dataset, we apply the knockoffs procedures that we are comparing. For each model, each signal amplitude, and each of the two methods, we measure performance in various different ways. First, for all the methods we estimate the power as
$\textnormal{Power}=\frac{1}{D}\sum_{d=1}^{D}\frac{|R_{d}\cap I_{1}|}{|I_{1}|},$
where $R_{d}$ denotes the rejected set in the $d$-th simulation run, and $I_{1}$ is the set of non-null (or relevant) variables.
Next, to assess Type I error control, we report different metrics depending on whether the procedure targets the FDR or the PFER.
5.2.1 For the methods that control the FDR (Sections 5.3.1, 5.3.3 and 5.3.2)
We measure the average false discovery proportion (i.e., empirical FDR) across the simulations:
$\textnormal{Average FDP}=\frac{1}{D}\sum_{d=1}^{D}\frac{|R_{d}\cap I_{0}|}{|R_{d}|\vee 1},$
@@
$\textnormal{Average ratio FDP/}{\alpha}=\frac{1}{D}\sum_{d=1}^{D}\frac{|R_{d}\cap I_{0}|}{|R_{d}|\vee 1}\frac{1}{{\alpha}_{d}},$
@@
5.2.2 For the methods that control the PFER (Section 5.3.4)
We measure the average number of false discoveries across the simulations:
$\textnormal{Average FD}=\frac{1}{D}\sum_{d=1}^{D}{|R_{d}\cap I_{0}|}.$
Furthermore, we consider the deviation of the ratio $\frac{FD{\eta}}{\nu}$ from $1$ to validate that eq. (12) is indeed satisfied and investigate the tightness of the control. For the derandomized knockoffs that control the PFER, following the configuration in Ren et al. (2023), we set $\eta=0.50$. For our post-hoc approach, the corresponding level is the data-dependent quantity $\tilde{\eta}^{\textnormal{ph}},$ returned by Algorithm 3. More precisely, we consider
$\textnormal{Average ratio FD}{\eta}/\nu=\frac{1}{100}\sum_{d=1}^{100}\frac{|R_{d}\cap I_{0}|{\eta}_{d}}{\nu},$
where ${\eta}_{d}$ is the parameter $\eta$ obtained in the $d$-th simulation run.
@@ 5.3 Results
This section is structured as follows. In Subsection 5.3.1, we begin by comparing our post-hoc knockoff procedure with the original knockoff filter of Candès et al. (2018). In Subsection 5.3.2, we then compare our method with the recently proposed calibrated knockoff procedure of Luo et al. (2025), which overcomes a key limitation of the original approach by mitigating the thresholding phenomenon. Next, in Subsection 5.3.3, we compare our derandomized post-hoc variant with the derandomized knockoffs of Ren and Barber (2024) designed for FDR control. Finally, in Subsection 5.3.4, we complete the analysis by evaluating our method against the derandomized knockoffs of Ren et al. (2023) for controlling PFER.
5.3.1 Comparing against original knockoffs (Candès et al., 2018)
@@
In this section we illustrate the performance of our proposed post-hoc knockoff (presented in Algorithm 1) and compare to the original knockoff procedure from Candès et al. (2018). The simulation setup is the low-dimensional regime presented in Section 5.1. Through the simulation we set the $\alpha^{\textnormal{kn}}$ level to $0.20$, which corresponds to a Type I error (FDR) of $0.20$ for the traditional knockoffs, and also serves as the initial significance level for our post-hoc knockoffs (Algorithm 1).
@@
In Figure 2, we evaluate the performance of the different methods as a function of $p_{\text{relevant}}$, with signal strength held fixed. When the number of truly relevant variables is small, traditional methods suffer from very low power. In contrast, post-hoc knockoffs offer the flexibility to substantially increase power by increasing the nominal level when the original method does not make any rejections. Across all settings, the average FDP remains close to the nominal target level $\alpha^{\textnormal{kn}}$, similar to that of the original knockoff procedure. Furthermore, when power is similar between the two methods (e.g., Gaussian setting for $p_{\text{relevant}}>6$), the post-hoc approach can achieve stricter error control, resulting in effective $\alpha$ values less than $0.20$. In the last row, we plot the average $\textnormal{FDP/}{\alpha}$ ratio, which remains consistently below 1. This confirms that the controlled error rate, as defined in equation (4), is indeed maintained.
@@
Similar behavior is observed when $p_{\text{relevant}}$ is held constant and the signal amplitude is varied, as shown in Figure 3. Post-hoc knockoffs consistently achieve higher power than traditional methods, especially at moderate signal strengths. For high signal amplitudes, our method achieves high power, and at the same time, it can provide more stringent control, i.e., $\alpha<0.20.$ Interestingly, the average FDP remains below the nominal $\alpha$ level in all scenarios. Furthermore, the average $\textnormal{FDP/}{\alpha}$ ratio remains below 1, again confirming that the error control in equation (4) holds.
Refer to captionRefer to captionRefer to caption
(a) Gaussian
Refer to captionRefer to captionRefer to caption
(b) Logistic
Figure 2: Performance comparison between the proposed derandomized post-hoc knockoff procedure and the original derandomized knockoffs method of Candès et al. (2018) across various numbers of actual relevant variables $p_{\text{relevant}}$. The signal amplitude is fixed at $A=8$ and $A=14$ for the Gaussian and Logistic model respectively. The first row reports the realized power, the second row reports the nominal level $\alpha$ (for the original method this is the user specified level $\alpha^{\textnormal{kn}}$, whereas for the post-hoc knockoff procedure it is the data-dependent level $\tilde{\alpha}^{\textnormal{ph}}$ returned by Algorithm 1) together with the estimated average FDP, and the last row reports the average ratio $\textnormal{FDP/}{\alpha}$. All values are averaged over 2000 runs.
Refer to captionRefer to captionRefer to caption
(a) Gaussian
Refer to captionRefer to captionRefer to caption
(b) Logistic
Figure 3: Performance comparison between the proposed derandomized post-hoc knockoff procedure and the original derandomized knockoffs method of Candès et al. (2018) across various signal amplitude values AA. The number of relevant variables is fixed at $p_{\text{relevant}}=6$. The first row reports the realized power, the second row reports the nominal level $\alpha$ (for the original method this is the user specified level $\alpha^{\textnormal{kn}}$, whereas for the post-hoc knockoff procedure it is the data-dependent level $\tilde{\alpha}^{\textnormal{ph}}$ returned by Algorithm 1) together with the estimated average FDP, and the last row reports the average ratio $\textnormal{FDP/}{\alpha}$. All values are averaged over 2000 runs.
Key takeaways from the comparison with the original knockoffs: Our post‑hoc derandomized knockoff procedure lead to power gains in weak‑signal regimes by increasing the significance level in cases where the original method does not make any discoveries. In addition, in scenarios where both methods attain similar power, the post‑hoc approach can still provide tighter error control, reflected in lower $\alpha$ values. It should be noted that our improvement is a “free-lunch”: It only reports a larger significance level if the original method did not produce any rejections.
5.3.2 Comparing against calibrated knockoffs (Luo et al., 2025)
In this section we illustrate the performance of our proposed post-hoc knockoff (presented in Algorithm 1) and compare to the calibration knockoffs method of Luo et al. (2025). Calibration knockoffs were proposed to address a limitation of the original knockoff filter: because the selection threshold is data-dependent, it can become overly conservative, leading to an empty or very small rejection set (the so-called thresholding phenomenon) even when the signal is not negligible. The calibration knockoffs approach mitigates this issue by calibrating the selection threshold using additional randomness, aiming to stabilize the cutoff and increase the probability of making discoveries while maintaining the desired Type I error control. The simulation setup is the low-dimensional regime presented in Section 5.1. Through the simulation we set the $\alpha^{\textnormal{kn}}$ level to $0.20$ FDR, i.e., this is the nominal level of the calibration knockoffs, and the initial significance level of our post-hoc knockoffs.
Figure 4 reports results under the Gaussian linear and logistic regression models with $p=50$, for two sparsity levels $p_{\text{relevant}}\in\{10,15\}$. Across all configurations, the two procedures exhibit very similar power curves as the signal amplitude increases. For weak signals (Gaussian: amplitudes $\leq 3$; Logistic: amplitudes $\leq 7$) the posthoc method leads to higher $\alpha$ levels of FDR, while for moderate to large signals (Gaussian: amplitudes >3>3; Logistic: amplitudes >7>7), the post-hoc approach typically yields more stringent error control, leading to smaller $\alpha$ levels. We emphasize that the estimated average FDP is very similar for the two methods and remains below the nominal level $0.20$ throughout. Furthermore, both methods ensure that the average ratio $\mathrm{FDP}/\alpha$ remains below 1.
Refer to captionRefer to captionRefer to caption
(a) Gaussian
Refer to captionRefer to captionRefer to caption
(b) Logistic
Figure 4: Performance comparison between the proposed post-hoc knockoff procedure and the calibration knockoffs method of Luo et al. (2025) under Gaussian and Logistic models with $p=50$. Results are reported across signal amplitudes for varying sparsity levels, given by the number of relevant variables $p_{\text{relevant}}=10$ and $15$. The first row reports the realized power, the second row reports the nominal level $\alpha$ (for the calibration method this is the user specified level $\alpha^{\textnormal{kn}}$, whereas for the post-hoc knockoff procedure it is the data-dependent level $\tilde{\alpha}^{\textnormal{ph}}$ returned by Algorithm 1) together with the estimated average FDP, and the last row reports the average ratio $\textnormal{FDP/}{\alpha}$. All values are averaged over 2000 runs.
Key takeaways from the comparison with the calibrated knockoffs: Across all scenarios, post‑hoc knockoffs achieve power nearly identical to that of calibrated knockoffs. With regard to the Type I error rate $\alpha,$ our post-hoc method leads to higher level for weak signals, but it also leads to stricter error control for moderate to large signals. Apart from this performance, our method is offering several practical and methodological advantages. First, our procedure is straightforward to apply: in contrast to calibrated knockoffs, which implement conditional calibration via per‑variable fallback tests, conditional‑null sampling, and budget accounting, our method requires neither calibration nor auxiliary randomness and does not rely on additional distributional assumptions. Hence, it is compatible with, and can be straightforwardly used with, any knockoff statistics, including nonparametric scores such as those derived from random‑forest importance measures. Second, in contrast to the calibrated knockoff method, the computational overhead of post‑hoc knockoffs is negligible, adding virtually no cost beyond running the original knockoff filter. Most importantly, the method admits a natural derandomization, and in the next two subsections we examine how this derandomized versions compare with other state‑of‑the‑art derandomized procedures.
5.3.3 Comparing against original derandomized knockoffs for controlling fdr (Ren and Barber, 2024)
@@
In this section, we illustrate the performance of our proposed derandomized post-hoc knockoff procedure for controlling the FDR (presented in Algorithm 2) and compare it with the original derandomized knockoff procedure from Ren and Barber (2024). The simulation setup is the high-dimensional regime presented in Section 5.1. For the nominal levels, we adopt the configurations from Ren and Barber (2024): we set the FDR level $\alpha=0.20$, which corresponds to the $\alpha_{\textnormal{ebh}}$ level of the derandomized knockoffs and serves as the initial significance level in our post-hoc derandomized knockoffs procedure (Algorithm 2), while in both methods we set $\alpha^{\textnormal{kn}}=\frac{\alpha_{\textnormal{ebh}}}{2},$ a configuration suggested by Ren and Barber (2024).
@@
Figure 5(a) presents the results under the Gaussian linear model, where the dimensionality is $p=800$. When $p_{\text{relevant}}=80$, which is the setting considered in Ren and Barber (2024), both methods exhibit comparable performance overall. However, our post-hoc procedure yields a modest power gain at lower signal amplitudes by increasing the nominal level in scenarios where the original derandomized method does not make any rejections. This becomes more pronounced in sparser settings, where $p_{\text{relevant}}=20$. Notably, in moderately sparse scenarios, our method can also achieve a lower average Type I error rate while maintaining higher power. For example, when $p_{\text{relevant}}=20$ and the signal amplitude exceeds $5$, our procedure attains similar power with improved Type I error control. The results are similar for the logistic model provided in Figure 5(b).
Refer to captionRefer to captionRefer to caption
(a) Gaussian
Refer to captionRefer to captionRefer to caption
(b) Logistic
@@
Figure 5: Performance comparison between the proposed derandomized post-hoc knockoff procedure for controlling FDR against the original derandomized knockoffs method by Ren and Barber (2024) under Gaussian ($p=800$) and Logistic models ($p=600$). Results are reported across signal amplitudes for varying sparsity levels, given by the number of relevant variables $p_{\text{relevant}}$, including the original setting: $p_{\text{relevant}}=80$ for the gaussian, and $p_{\text{relevant}}=60$ for the logistic. The first row reports the realized power, the second row reports the nominal level $\alpha$ (for the calibration method this is the user specified level $\alpha^{\textnormal{ebh}}$, whereas for the post-hoc knockoff procedure it is the data-dependent level $\tilde{\alpha}^{\textnormal{dph}}$ returned by Algorithm 2) together with the estimated average FDP, and the last row reports the average ratio $\textnormal{FDP/}{\alpha}$. All values are averaged over 200 runs.
Key takeaways from the comparison with the derandomized knockoffs for controlling FDR: Our post‑hoc derandomized knockoff procedure yields meaningful power improvements in weak‑signal settings by increasing the nominal level in scenarios where the original method does not make any discoveries. Furthermore, in scenarios where both methods achieve comparable power, the post‑hoc approach can yield tighter error control. As described before, our improvement can be seen as a “free-lunch”.
5.3.4 Comparing against original derandomized knockoffs for controlling pfer (Ren et al., 2023)
In this section, we illustrate the performance of our proposed derandomized post-hoc knockoff procedure for controlling the PFER (presented in Algorithm 3) and compare it with the original derandomized knockoffs procedure of Ren et al. (2023). The simulation setup is the low-dimensional regime presented in Section 5.1. For the nominal levels, we set the expected number of false discoveries $\nu=1$, and, following the configurations from Ren et al. (2023), we set $\eta=0.50$. In the original derandomized PFER knockoffs procedure, the parameter $\eta\in(0,1]$ is a prespecified threshold that determines how consistently a variable must be selected across the $k$ knockoff runs to be included in the final rejection set $R_{\eta}^{\textnormal{RWC}}$ in (11); thus, smaller $\eta$ makes rejections easier and larger $\eta$ makes them more stringent. Ren et al. (2023) showed that RWC-knockoff controls the PFER at level $\nu/\eta$, i.e., in our setting this is equal to $1/0.50=2$. The parameter $\eta$ is not treated as a fixed design parameter in our post-hoc derandomized knockoffs framework: instead, it is adjusted in a data-dependent way, as permitted by the guarantee in Theorem 7.
Figure 6 reports results under the Gaussian linear and logistic regression models with $p=50$, for two sparsity levels $p_{\text{relevant}}\in\{5,10\}$. Across all configurations, the two procedures exhibit very similar power curves as the signal amplitude increases.
Refer to captionRefer to captionRefer to caption
(a) Gaussian
Refer to captionRefer to captionRefer to caption
(b) Logistic
@@
Figure 6: Performance comparison between the proposed derandomized post-hoc knockoff procedure for controlling PFER against the original derandomized knockoffs method by Ren et al. (2023) under Gaussian and Logistic models with $p=50$. Results are reported across signal amplitudes for varying sparsity levels, given by the number of relevant variables $p_{\text{relevant}}=5$ and $10.$ The first row reports the estimated power, the second row reports the data-dependent level $\alpha$ together with the estimated FD, and the last row reports the average ratio $\textnormal{FD}*\eta/\nu$. All values are averaged over 2000 runs.
Interestingly, for moderate to large signals (Gaussian: amplitudes >3>3; Logistic: amplitudes >7>7), the post-hoc approach typically yields more stringent error control, leading to smaller estimated values of the expected number of false discoveries (generally below $2$). Finally, across all comparisons the ratio $FD*\eta/\nu$ for both methods remains bellow $1$, confirming that PFER is appropriately controlled.
Key takeaways from the comparison with the derandomized knockoffs for controlling PFER: The post‑hoc derandomized knockoffs achieve power similar to the original derandomized PFER method, while their adaptive thresholding enables more stringent error control whenever the signal strength is sufficiently high. It should be noted that our post-hoc implementation for PFER control does not always lead to larger power than the original method. However, if the $(R_{\eta}^{\textnormal{RWC}},\eta)$ combination of the original method seems more desirable, we could also switch to that combination post-hoc.
6 Applications in analyzing clinical trial data
In a recent work, Zimmermann et al. (2024) illustrate how knockoff-based methods can be applied to clinical trial data to identify prognostic and predictive biomarkers while controlling type-I errors in exploratory variable selection. In this section we will present applications of our post-hoc knockoffs methods in analysing clinical trial data: Subsection 6.1 will focus on the identification of prognostic variables using a synthetic dataset designed to mimic a real randomized controlled trial, whereas Subsection 6.2 will analyse predictive biomarkers using actual clinical trial data from five phase III studies.
6.1 Derive prognostic variables using post-hoc knockoffs
@@
To demonstrate the practical utility of our approach and the advantages of post-hoc knockoffs for analysing clinical trial data, we consider a synthetic dataset designed to mimic a real randomized controlled trial, introduced by Sun et al. (2024). To generate the data we used benchtm R package (Sun et al., 2024), using Scenario 1 with a continuous outcome and a sample size of $n=1000$. Each patient is described by $p=30$ baseline covariates, a dimensionality that reflects realistic clinical settings where tens of demographic, clinical, laboratory, and lifestyle variables are routinely collected and jointly analysed. The outcome for each patient was generated as a function of covariates $\mathbf{X}=(X_{1},X_{2},\dots,X_{30})$ and the binary treatment assignment $T\in\{0,1\}$ according to
@@
where $\mathbb{I}(\cdot)$ denotes the indicator function, $s=2.32$ controls the baseline covariate effect, and $\beta_{0}=0.053$ and $\beta_{1}=0.383$ determine the average treatment effect and the additional effect in the specified subpopulation, respectively. These parameters were tuned to mimic real clinical trial data, and a full description of the data-generating mechanism and simulation design is provided in Sun et al. (2024).
We simulate 10 independent RCT datasets following the procedure described above, denoted by R1, …, R10. To generate the knockoffs we used the sequential knockoff procedure (Kormaksson et al., 2021), implemented in the knockofftools package (Zimmermann et al., 2024). Then, we use the Lasso Coefficient Difference (LCD) knockoff statistic, which is computed as the difference between the absolute LASSO coefficient of each original variable and that of its knockoff copy, as the knockoff filter. Finally, we examine which variables are selected by the two methods (original knockoffs vs post-hoc knockoffs) under a nominal FDR level of 0.20. Figure 7 displays the corresponding heatmaps of selected variables. We remind the reader that, according to the data-generating process specified in equation (13), the set of truly relevant variables in this scenario is $I_{1}=\{T,X_{1},X_{11}\}.$
@@
As shown in the left panel, the original knockoffs method selects variables only in two runs out of 10, in the RCT 1 and RCT 9. If we use this method in practice, then most of the time the method makes no discoveries, leading to frequent failures in detecting true signals. In contrast, the post‑hoc method (right panel) selects variables $X_{1}$ and $X_{11}$, both true predictors, in all ten runs, and additionally selects the third relevant variable $T$ in 8 out of 10 runs. This indicates that, in practical applications, the method would almost always recover the truly relevant variables. While the average $\alpha^{\textnormal{ph}}$ across the ten runs is 0.3, it is worth emphasizing again that these gains come at no cost, since our method is more informative. In the two runs (RCT 1 and 9) where the original method makes discoveries, our post-hoc procedure identifies the same variable, with nominal levels of $0.20$ and $0.14,$ which is smaller than the nominal level used by the original KO. In the remaining eight runs, where the original KO makes no discoveries, our method provides more informative results: we almost always recover the true signals, and the average nominal $\alpha^{\textnormal{ph}}$ level across these runs is 0.33.
6.2 Derive effect modifiers in real trial
Identifying effect modifiers, also known as predictive biomarkers, is a central task in the analysis of clinical trial data. Effect modifiers are baseline patient characteristics that influence the magnitude or direction of a treatment effect. Detecting such variables is critical for a range of scientific and clinical activities, including defining subgroups with enhanced treatment benefit, informing treatment-personalization strategies, and generating hypotheses for further biological or mechanistic investigation. In this section study, we focus on the task of identifying effect modifiers from clinical trial data using our knockoffs and controlling FDR.
We will analyse five Phase III clinical trials of secukinumab (Cosentyx) for psoriatic arthritis (PsA): FUTURE 1–5 (Mease et al., 2015; McInnes et al., 2015; Nash et al., 2018; Kivitz et al., 2019; Mease et al., 2018). The combined dataset contains $n=1937$ patients across the five trials. Covariates with more than 20% missingness were removed. Remaining variables were imputed using multiple imputation, and completed datasets were aggregated using medians (numeric variables) or modes (categorical variables). This produced a final set of 70 mixed-type baseline variables including demographics, disease history, laboratory biomarkers, quality-of-life assessments, and the proteomic marker BD-2.(Cardner et al., 2023) Further details on how we generated the analysis dataset are provided in Sechidis et al. (2025). The primary endpoint is the binary ACR50 response at week 16 (Felson et al., 1993) and treatment heterogeneity is evaluated on the risk-difference scale, $\Pr(Y=1\mid T=1)-\Pr(Y=1\mid T=0).$
Building on our previous work by Sechidis et al. (2025), we first derive pseudo-observations of the individual treatment effect using the doubly robust (DR) learner framework. These pseudo-observations summarize treatment heterogeneity at the patient level in a model-agnostic manner. In that earlier analysis, the authors reported rankings of variable‐importance scores that were obtained using conditional inference trees and permutation importance metrics (Sechidis et al., 2025, Figure 17). The top ten variables in that ranking were C-reactive protein, age, proteomic marker BD-2, baseline fatigue score, apolipoprotein A1, alkaline phosphatase, total protein, sex, lymphocytes/leukocytes percentage, and ethnicity. Several of these variables, most notably C-reactive protein, Age, Fatigue Score, and proteomic marker BD-2, correspond to known effect modifiers previously identified in independent analyses (Sechidis et al., 2021; Bornkamp et al., 2024; Cardner et al., 2023).
While such rankings provide valuable descriptive insight, they do not by themselves deliver statistically controlled variable selection. In exploratory subgroup discovery and predictive biomarker research, methods such as knockoffs are essential for producing reproducible findings under explicit FDR control. To generate the knockoffs and compute the corresponding knockoff statistics, we followed the same procedure as in the previous section, using sequential knockoffs and the LCD statistic.
Firstly, we explore the selections obtained from the original derandomized knockoff filter with those from our proposed post-hoc knockoff procedure. For the nominal levels, we adopt the configurations from Ren and Barber (2024): we set the FDR level $\alpha=0.20$, which corresponds to the $\alpha_{\textnormal{ebh}}$ level of the derandomized knockoffs and serves as the initial significance level in our post-hoc derandomized knockoffs procedure, while in both methods we set $\alpha^{\textnormal{kn}}=\frac{\alpha_{\textnormal{ebh}}}{2}.$
Original derandomized knockoffs.
At the prespecified nominal FDR level of $0.20,$ corresponding to the $\alpha_{\text{EBH}}$ threshold of the derandomized knockoff filter, the original derandomized knockoff procedure did not select any variable, indicating that no covariate exhibited sufficiently strong evidence to be flagged as a potential effect modifier.
Post-hoc derandomized knockoffs.
At the post-hoc adjusted significance level $\alpha\approx 0.27$, the derandomized knockoff procedure selected four variables as effect modifiers. Thus, with a small relaxation of the target FDR (from $20\%$ to $27\%$), the post-hoc adjustment yielded a non-empty and interpretable set of promising effect modifiers.
Having established the contrasting selection behavior of the traditional and post-hoc derandomized knockoff procedures, we next examine how these selections relate to the underlying signal structure captured by the post-hoc derandomized knockoff procedure. Figure 8 presents the variable-importance ranking, computed as the average (compound) e-values across the derandomized runs. This ranking offers a transparent view of the covariates exhibiting the strongest evidence of effect modification. The four variables selected by our post-hoc derandomized knockoff method at the relaxed FDR level of $0.27$ are highlighted in red, and they are: C-reactive protein, baseline fatigue score, age and a proteomic marker BD-2. These selections closely mirror findings reported in prior analyses of these datasets, where the same variables were highlighted as leading drivers of treatment heterogeneity in PsA (Sechidis et al., 2021; Bornkamp et al., 2024; Cardner et al., 2023).
Figure 8: Variable importance ranking returned from the post-hoc derandomized knockoffs. This figure shows the top variables ranked by their importance measured as the average of (compound) e-values across the derandomized runs. Higher-ranked variables indicate stronger modification effects. We have highlighted with red the variables that got selected by our post-hoc derandomized knockoff procedure, and that control the FDR to the level of $0.27.$
As we saw in Section 4.1, the e-value framework allows us to move beyond a fixed rejection set and instead evaluate the specific level of evidence for any subset of interest. To illustrate this flexibility, consider a scenario where a practitioner is interested specifically in the evidence supporting only Age and the proteomic marker. Utilizing the property derived in Proposition 6, we can determine that the smallest nominal level $\alpha$ required to support this specific rejection set is approximately $0.52\%$. While our method enables this capability of determining the error level after choosing the rejection set, further research is needed to fully understand the practical consequences and the specific clinical contexts where such post-hoc decisions are most appropriate.
To sum it up, our post-hoc derandomized knockoff approach recovers clinically plausible modifiers while maintaining explicit FDR control, demonstrating that a modest relaxation of the nominal level can substantially improve power without inflating spurious discoveries. Together, these results illustrate that post-hoc adjustment offers a principled and practical enhancement to derandomized knockoffs, enabling robust and interpretable identification of effect modifiers in high-dimensional clinical trial settings.